Suppose your top drawer contains different colored socks: 4 are white, 16 are black, 6 are pink, and 12 are blue. All socks in the drawer are loose (unpaired). In the morning, you randomly select two socks, one at a time. Calculate the following probabilities, writing your answer either as a decimal or a fraction.

(a) What is the probability that you get a blue pair of socks?



(b) What is the probability that you do not get a blue pair of socks?



(c) What is the probability that you get either a white pair or a blue pair of socks?



(d) What is the probability that you get one black sock and one white sock?
Answer #1

a)

P( get a blue pair of socks )=(12/38)*11/37)=66/703

b)

P( do not get a blue pair of socks )=1-66/703 =637/703

c)

probability that you get either a white pair or a blue pair of socks

=(4*3+12*11)/(38*37)=72/703

d)

probability that you get one black sock and one white sock =2*4*16/(38*37)=64/703